OpenStudy (anonymous):

Hi, Have a completing the square integral I am trying to do and need some help getting the completing the square part down.

OpenStudy (anonymous):

\[\int\limits (x^2-2x+1)/\sqrt{(x^2-2x+10)}\]

OpenStudy (anonymous):

alright well do you know how to complete a square?

OpenStudy (anonymous):

So far I have pulled Moved 10 over under the radical and replaced it with 1 so that half of the middle term squared would be the third term. Im not sure where to subtract one to make up for it though......

OpenStudy (anonymous):

i think your supposed to subtract one from the other 1.

OpenStudy (anonymous):

So my new equation will look like \[\int\limits (x^2-2x+1)\div \sqrt{9+(x-1)^2}\]

OpenStudy (anonymous):

then a tan sub?

OpenStudy (anonymous):

YES YOUR EQUATION IS CORRECT BUT YOU NEED TO BE CAREFULLY WHEN YOU DO THE EQUATION.

OpenStudy (anonymous):

I think I am almost done.....Im at the integral of sec^3theta-sectheta

OpenStudy (anonymous):

OK WHEN YOU HAVE THE ANSWER JUST TEEL ME

OpenStudy (anonymous):

I could use help with that though (;

OpenStudy (anonymous):

Do I use the power reducing identity for sec^3(theta)?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

DO YOU GET THE ANSWER JANSON

OpenStudy (anonymous):

not yet......trying to find the identity.

OpenStudy (anonymous):

I WILL BE HERE IF YOU HAVE ANY QUESTION.

OpenStudy (anonymous):

Are you sure I dont do sec^3 with integration by parts?

OpenStudy (anonymous):

WELL YOU CAN DO IT BUT IT WIL HAVE MORE WORK TO DO.

OpenStudy (tkhunny):

What are you folks doing? Notice the similarity of the numerator. \(\int\limits \dfrac{x^{2}-2x+1}{\sqrt{x^2 - 2x + 10}}\;dx = \int\limits \dfrac{(x-1)^2}{\sqrt{(x-1)^2 + 9}}\;dx\) I found the substitution \(u^{2} = (x-1)^{2} + 9\) quite useful.

OpenStudy (anonymous):

I KNOW THE THE SIMILARITY OF THE NUMERATOR. I JUST TRYING TO HELP HIM

OpenStudy (anonymous):

So I found the identity but right now my integral is \[9intSec^3\theta d\] So when I use the power reducing indentify do I pull the constant 9 out in front of it also? The point of the problem for me on this one is to complete the square and then use trig sub

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Can you help walk me through some of this last part?

OpenStudy (anonymous):

SHOW ME THE EQUATION

OpenStudy (anonymous):

\[9\int\limits \sec^3\theta-\sec \theta d \theta \]

OpenStudy (anonymous):

CAN YOU HURRY UP BECAUSE I HAVE TO DO MY HOMEWORK FOR AP CHEMISTRY AND AP US HISTORY

OpenStudy (anonymous):

its right above you...

OpenStudy (anonymous):

As in would you show me how to do this part.

OpenStudy (anonymous):

you dont seem to know what your doing );

OpenStudy (anonymous):

This is Cal 3

OpenStudy (anonymous):

its wrong.

OpenStudy (anonymous):

IS NOT BECAUSE MY BROTHER IS HELPING ME MY FRIEND